OpenStudy (anonymous):

convert y = x^2 -2x -3 into vertex form

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

Ok, well first you take the coefficient of x, divide it by 2, then square the result. so, (-2/2)^2 -1^2 1 you add and subtract that into the equation y=x^2-2x+1-1-3 Factor the first part which is now a perfect square y=(x-1)^2-1-3 simplify y=(x-1)^2-4 understand how i got that?

OpenStudy (anonymous):

yea